OpenStudy (anonymous):

what the relative uncertainity of an object of length 2 cm what is the the relative uncertainity is we use the above meter

OpenStudy (anonymous):

meter stick = o.1cm vernier = 0.01 cm micrometer = 0.001cm relative uncertainity for 2cm will be for meter stick = (0.1/2 )*100% = 5% vernier = (0.01/2)*100% = 0.5% micrometer = (0.001/ 2)*100% = 0.05%

OpenStudy (anonymous):

dont i have to convert cm to mm

OpenStudy (anonymous):

for the 2

OpenStudy (anonymous):

no i had already worked the precisions in cm

OpenStudy (anonymous):

another person told me if you measure it to be 2cm with any of the above instruments then the measurement will look like this 2cm±precision or 2cm±uncertainty so report plus or minus your precision as your uncertainty.

OpenStudy (anonymous):

yea thatz true

OpenStudy (anonymous):

so how do i use what this person said

OpenStudy (anonymous):

so 4 meter stick= 2cm± 0.05 vernier = 2cm± 0.005 miocrometer= 2cm± 0.0005 this means that measurements using a micrometer are more precise and accurate

OpenStudy (anonymous):

what i had given before was the % precisions using the various instruments.

OpenStudy (anonymous):

what is the precision of these instruments